• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name slifer

    Log entry time 18:26:37 on October 26,1998

    Entry number 3752

    keyword=Bscope

    17:55 Bscope has been performed by MCC.
    The range is still only to 3kHz,
    but a quick look at their elog shows
    that Alex's 30kHz request caused a slew
    of software problems. Contacted MCC and was told
    that they are waiting for morning when their software
    people return to fix this.
    (Note: It would have been a good idea to tell the shift workers what
    Bscope was useful for before now instead of just instructing them
    to request it on each shift. The problem would have been caught
    much sooner.)